# Which Similarity-Sensitive Entropy (S-entropy)?

### v3

@@ 4.3 Bounds on LCR and VS
When $Z$ is full rank but otherwise as in 1, for all $q\in[-\infty,0]$
$VS_{q}(Z)\geq n\geq D_{q}(Z,\frac{1}{n})$ (10)
@@
$D_{-\infty}(Z,\frac{1}{n})\geq D_{q}(Z,\frac{1}{n})$ (11)
$D_{-\infty}(Z,\frac{1}{n})=\frac{n}{\displaystyle\min r_{i}}\leq n$ (12)
A computational search of $\approx 4,000$ examples involving similarity matrices of up to size $50\times 50$ failed to find counterexamples for any of several values of $q$, including $q=0$ and $q=1$. The method behind this search has been successfully used to found counterexamples to other inequalities [8], adding further support for Conjecture 1 being true. We present an approach to a proof for all $q$ in Appendix A.3 and present an additional proof on bounds in Appendix A.4.
5 Discussion
Quantifying the information contained in a system is essential for virtually any complex system. This includes ML datasets, where the fact that most elements are generally unique—no repeated images (in image datasets) or rows (in tabular datasets)—renders traditional entropy (Shannon entropy and the other Rényi entropies) uninformative. This is because traditional entropy measures information encoded by the shape of the elements’ frequency distribution, but when the elements each appear only once, this distribution is flat (Fig. 1): there is no additional information there. Traditional entropy is used to measure the distribution of class sizes, which often vary; this is class balance. However, there is no obvious way to use class balance to measure the the information contained within or between classes, which is a property of the elements themselves.
5.1 Richer measures of the information that ML datasets contain
Conceptually, the main difference between clustering and S-entropy is that S-entropy accounts for similarities both within and between clusters, whereas most clustering algorithms (and classifiers) lack this ability: effectively, the within-cluster (or intra-class) similarity is always 1 and the between-cluster similarity is always 0. The exception is fuzzy clustering, in which each element can belong to more than one cluster, with each element being parameterized by its strength of membership for one or more clusters. One can view LCR as carrying fuzzy-clustering to its logical limit, with each unique element being its own cluster and the similarities giving other elements’ strength of membership in that cluster and thereby contributing to the size of the cluster. This is a fuzzy-clustering interpretation of the ordinariness, $Z\mathbf{p}_{i}$: a big cluster to which many unique elements belong fairly strongly would be considered “ordinary” (Fig. 2c). This perspective provides another way of seeing that LCR is the entropy of a system after accounting/adjusting/correcting for the similarity among its elements.
How do LCR and VS compare with the number of clusters identified by HDBSCAN? Predicting the relationship is difficult: large clusters whose peripheral elements are relatively distinct tend to increase the LCR/VS values, whereas strong similarities among elements—both within a cluster and across neighboring clusters—tend to reduce them (see Figs. 1, 5c). Indeed, many of the datasets we examined exhibited a pronounced difference in magnitude between tabular and imaging datasets when comparing the HDBSCAN cluster count with LCR and VS. For tabular data sets, both LCR and VS almost always surpassed the HDBSCAN cluster count, often by 1-2 orders of magnitude, reflecting that many rows remain unclustered by hard clustering (these do not contribute the the HDBSCAN total).
In contrast, for imaging datasets, LCR and VS were typically 1-2 orders of magnitude lower than the number of clusters, indicating substantial similarity among elements across clusters, which hard clustering does not capture (Fig. 2c). We interpret these systematic differences as evidence of the advantages of S-entropy relative to hard clustering, for describing a system’s structure. We do not regard this as a failure of HDBSCAN; a limitation of our study that we did not explore a range of HDBSCAN hyperparameters, since the goal was simply to illustrate the general relationship between similarity‑sensitive entropy and clustering. However, we note that by the “eye test,” UMAP plots validate HDBSCAN’s clustering, suggesting hyperparameter search is unlikely to meaningfully affect the findings or conclusions presented.
5.2 The effect of scaling the similarity measure/similarity matrix
We demonstrated that the relationship between LCR and VS depends strongly on how the similarity‑ matrix values are scaled, which we illustrated by introducing the half‑distance parameter $k$. Recall that similarity between two elements ranges from 0 to 1; scaling modifies this relationship by raising the values to a power, thereby shifting the curve in the middle of this range, while maintaining the anchoring at each end (at 0 and 1).
Why should one need to scale the similarity between elements in the first place? Should scaling be permitted at all? When similarity is derived from external knowledge or some principle or outside requirement, we argue that scaling should not be applied: the external knowledge defines and thereby constrains the similarity function a priori, its output for any given pair is fixed, and the similarity matrix simply collects these values for all pairs. An example is binding similarity in immune repertoires, defined as the relative dissociation constants of two antibodies (or TCRs) for a given antigen (or set of antigens) [2]: this is the unique definition that preserves biophysical additivity; arbitrary scaling would break this relationship and thereby produce a similarity matrix that is objectively incorrect.
In contrast, when similarity is heuristic and not constrained by such knowledge, as with the Euclidean‑distance-based similarity measures used here, there is no a priori reason to privilege the default similarity ($k=1$) over other values of $k$. For instance, RMSD provides a perfectly valid similarity definition and corresponds to Euclidean distance with $k=1/\sqrt{d}$, where $d$ is the dimension of the underlying data (number of columns for tabular data, number of pixels $\times$ number of color channels for image data, 2 for the umap embedding, etc). Therefore, in heuristic cases, scaling can serve as a reasonable dial that investigators can adjust to obtain values that make the most sense given the context/are maximally informative about the system. Such choices arise often in measurement. For example, the parameter $q$ is an analogous dial in traditional entropy (as well as S-entropy), allowing investigators to adjust the emphasis on frequency. As with $k$, prior information can constrain choice of $q$; for example, when the measurement has to do with collision frequency, $q=2$ should be used, consistent with $q=2$ corresponding to Simpson’s index. Whether specific values of $k$ might similarly be linked to known quantities of interest is left for future work.
A key feature of scaling is that it interpolates similarity‑sensitive entropy between 1 (for small $k$) and traditional entropy (for large $k$). When $k$ is small, the off‑diagonal entries of the similarity matrix approach 1, implying that all unique elements are mutually similar; in the limit the system behaves as if it contained a single unique element. Conversely, for large $k$ the off‑diagonal entries approach zero, yielding the identity matrix in the limit; the unique elements become completely dissimilar, and traditional entropy is recovered. In a system where the $n$ elements are all unique—typically the case for ML data sets, including the imaging and tabular data sets examined here—the effective‑number form of traditional entropy equals the dataset size. Consequently, for such uniform systems $k$ interpolates both LCR and VS from 1 to $n$. Other scaling approaches may exhibit the same property.
For heuristic similarity functions, which $k$ should one choose? Qualitatively, the answer is a value roughly midway between the extremes (Fig. 6). However, our quantitative findings indicate that the optimal choice is not obvious. We found that the correlation between LCR and VS is minimized at $k=0.2$ for tabular data sets and $3.4$ for imaging data sets—not at any obvious definition of the midpoint of the range (defined for example as the geometric mean of the $k$ values where the correlation first falls to 0.99 on each side). Moreover, the shapes of these curves, with their multiple extrema, do not appear to follow a simple parametric form, making it difficult to predict precisely where the two metrics are maximally mutually informative. We also observed that even at the $k$ that minimizes the LCR-VS correlation, the $R^{2}$ between LCR and VS remained relatively high in absolute terms, at $0.68$. Perhaps this was to be expected, since both metrics quantify the effective number of elements using the same underlying similarity measure, albeit in different ways (see Section 6). The dependence on $k$, however, was quite strong.
Our results suggest first testing whether the entropy lies near 1 or $n$; if so, to explore, if possible, different $k$ values to delineate a non‑trivial range; and finally to select a $k$ approximately in the middle of this range (on a logarithmic scale). Both the choice of similarity measure and the value of $k$ have effects on the resulting entropy that cannot be ignored. Thus to some extent, from a quantitative perspective, the information in a system is in the eye of the beholder.
6 Conclusion
So: LCR or VS? Our results support defaulting to LCR when the goal is to measure the similarity‑adjusted entropy of a system, and to use VS only in one of two special cases: (i) when the elements are usefully thought of as linear combinations of a set of mutually orthogonal ur‑elements/eigen‑elements, or (ii) when the system possesses a quantum‑mechanical character (see below). If the aim is a more general multidimensional characterization, there is no clear reason to avoid employing both metrics; keeping in mind that they are likely to correlate, our findings recommend selecting $k$ within the non‑trivial range (see Section 5.2) to maximize their independence and thereby their combined informativeness.
We note that, quantitatively, VS will be larger than LCR for any given $k$ and $q$. We observed this ordering empirically and proved it for a broad class of cases, outlining a proof for the remaining situations; a computational search uncovered no counter‑examples. This relationship can be rationalized by recognizing that, although VS is a similarity‑sensitive entropy, its computation includes a traditional entropy component: after incorporating similarity to derive the eigen‑elements, VS is the traditional entropy of the resulting eigenvalues; traditional entropy is always greater than or equal to similarity‑sensitive entropy.
From a practical perspective, LCR has three advantages. First, it is less computationally expensive whenever elements appear multiple times (i.e., when $\mathbf{p}$ is not flat/uniform), because the similarity matrix for LCR has order equal to the number of unique elements, whereas VS requires a matrix whose order equals the total number of elements; both calculations scale as the square of the matrix order. Empirically, we observed a speed advantage of sentropy over vendi-score even when all elements were unique, a situation that frequently occurs in ML datasets (Fig. 8).
Second, LCR does not require the similarity matrix to be PSD, whereas VS does. Similarity matrices need not be PSD; a perfectly reasonable set of similarity measurements can produce a matrix that fails the PSD condition, which would complicate the use of VS. Even symmetric similarity matrices—the most common and most natural type, satisfying $z_{ij}=z_{ji}$, so that similarity is reciprocal—are not guaranteed to be PSD, even when their entries obey the triangle inequality. The prevalence of such non‑PSD cases in practice is unknown; we note that all similarity matrices examined for our datasets were PSD (a consequence of how we chose our similarity measures; this follows from Theorem 2.11 in [16]). Thus, PSD‑ness is irrelevant for LCR but must be checked to apply VS.
And third, LCR in its various forms [3, 4] (Table 1) is part of an extensively characterized framework [17] that provides similarity‑sensitive analogs of mutual information, relative entropy, and cross‑entropy and offers methods for partitioning the overall similarity‑sensitive entropy ($\gamma$) into within‑class ($\alpha$) and between‑class ($\beta$) components. In entropic terms, $\alpha$ is the (effective‑number form of the) joint entropy of the elements together with their class label; $\beta$ is the mutual information between the elements and the class label; and $\gamma$ is the marginal entropy of the elements irrespective of class. In this work we computed LCR in its $\gamma$ form; when only a single class is present, $\gamma$ coincides with $\alpha$. Measuring the full $\alpha$, $\beta$, and $\gamma$ decomposition across ML classes is a promising direction for future research. (The sentropy package computes $\alpha$, $\beta$, and $\gamma$.)
The necessary mathematics for extending this decomposition to VS exists, because VS is based on von Neumann (quantum) entropy, for which quantum mutual information, quantum relative entropy, and related quantum quantities are well understood [24]. However, computing these quantities requires a matrix that simultaneously encodes each element and its class membership. If there are $n$ total elements and $k$ classes, the VS decomposition demands either an $nk\times nk$ matrix or an $n\times k\times n\times k$ 4-dimensional tensor (which can be reduced to $n\times n$ and $k\times k$ matrices by taking a partial trace). In contrast, LCR requires only the similarity matrix $Z$. Consequently, VS needs an explicit notion of similarity both between classes and between elements, whereas in LCR any class‑level similarity arises purely from element‑level similarities within the classes. In principle, the VS formulation would permit the confusing situation in which two classes contain exactly the same individuals but nonetheless have zero similarity—a scenario that cannot occur under LCR, in which the similarity between classes is purely a function of similarities among the elements they contain.
In sum, LCR and VS provide complementary descriptions of similarity‑sensitive entropy, a unifying concept that is likely to find applications in many domains in which traditional entropy is currently a mainstay. We anticipate that the deeper understanding of their relationship and properties will aid such investigations.
7 Acknowledgements
Appendix A Appendix
A.1 Python code for finding counterexamples to $VS\geq LCR$
@@
    evals = np.array([e for e in evals if e>0]) # keep only the support
@@
    return vendi_score(Z, q) - uniform_div(Z, q)
@@
    # map from reals to positive reals, reshape
@@
combos = [{’q’: q, ’n’: n, ’r’: r, ’rep’: rep, ’scale’: s} \
for q in qlist for n in sizes for s in scales \
for rep in replicates \
@@
    rows = pool.map(map_func, combos[ix_start:ix_end])
@@
    results_df.to_pickle(’../data/vendi_inequality_counterexample_search.pkl’)
A.2 Proof that $VS\geq LCR$ for $q\in\{2,3,\infty\}$
Theorem 1.
Let $Z$, $\text{VS}_{q}(Z)$, and $D_{q}(Z,\frac{1}{n})$ be as in 1. Then when $q\in\{2,3,\infty\}$,
$VS_{q}(Z)\geq D_{q}(Z,\frac{1}{n})$ (13)
Proof.
At $q=2$,
[…]
$\frac{n^{2}}{\sum_{i,j}Z^{2}_{i,j}}\geq\frac{n^{2}}{\sum_{i,j}Z_{i,j}}$ (15)
For $q=3$, we have that
[…]
This gives
[…]
proving the $q=3$ case.
For $q=\infty$, we invoke the Gershgorin circle inequality. We first note that
$\mathrm{VS}_{\infty}(Z)$ $=$ $\frac{n}{\lambda_{\text{max}}}$ (19)
$D_{\infty}(Z,\frac{1}{n})$ $=$ $\frac{n}{r_{\text{max}}}$ (20)
where $\lambda_{\text{max}}$ is the largest eigenvalue of $Z$, i.e. $\lambda_{\text{max}}=\displaystyle\max_{i}\lambda_{i}$, and likewise $r_{\text{max}}=\displaystyle\max_{i}r_{i}$. However, by the Gershgorin circle inequality, every eigenvalue of $Z$ must fall within a Gershgorin interval, i.e. in one of the intervals $Z_{ii}-\sum_{j\neq i}Z_{i,j},Z_{ii}+\sum_{j\neq i}Z_{i,j}]$. As such, noting that $Z_{ii}+\sum_{j\neq i}Z_{i,j}=\sum_{j}Z_{i,j}=r_{i}$, we must have that every eigenvalue $\lambda_{\mu}$ of $Z$ is bounded by $\lambda_{\mu}\leq r_{\text{max}}$. But this means that
A.3 Approach to a proof of $VS\geq LCR$ for all $q$
A proof for 1 at all $q$ except for $1$ (which could perhaps be filled in using continuity considerations) would follow from the following:
Conjecture 2.
Let $Z$ be as in 1, and denote by $\mathbf{1}$ the $n$ dimensional vector each of whose entries are $1$. Then for $q>1$,
[…]
Likewise, for $q<1$,
[…]
At $q>1$, 2 would in turn be proven by the following conjecture:
Conjecture 3.
Let $M$ be any $n\times n$ symmetric matrix whose entries obey $0\leq M_{ij}\leq 1$, and denote by $\mathbf{1}$ the $n$ dimensional vector each of whose entries are $1$. Then for any $\alpha\geq 0$
$\mathbf{1}^{\intercal}M^{\alpha}\mathbf{1}\leq\mathbf{1}^{\intercal}\left(M\mathbf{1}\right)^{\alpha}$ (24)
A.4 Other bounds
Beyond the conjectures upper bound of uniform diversity-with-similarity by Vendi score, we are able to prove the following bound in the opposite direction:
Theorem 2.
Let $Z$, $\text{VS}_{q}(Z)$, and $D_{q}(Z,\frac{1}{n})$ be as in 1. Then
Proof.
We begin by constructing the diagonal matrix $W$ according to
$W_{ii}$ $=$ $\frac{1}{T}\sum_{j}Z_{ij}$ (26)
$T$ $=$ $\sum_{i,j}Z_{ij}$ (27)
[…]
$\Rightarrow$
[…]
However, since $W$ is diagonal we have that
[…]
$=$ $-\sum_{i}\frac{1}{n}\log\frac{\sum_{j}Z_{ij}}{T}$ (31)
As such, we have that
$\mathrm{VS}_{1}(Z)\leq\frac{T}{n}D_{1}(Z,\frac{1}{n})$ (35)
and so the theorem is proved.

### v4

@@ Appendix B Other mathematical results
Let $Z$, $\text{VS}_{q}(Z)$, and $D_{q}(Z,\frac{1}{n})$ be as in 1. Then
@@
We begin by constructing the diagonal matrix $W$ according to
$W_{ii}$ $=\frac{1}{T}\sum_{j}Z_{ij}$ (48)
$T$ $=\sum_{i,j}Z_{ij}$ (49)
@@
$\Rightarrow$
[…]
However, since $W$ is diagonal we have that
[…]
$=-\sum_{i}\frac{1}{n}\log\frac{\sum_{j}Z_{ij}}{T}$ (53)
As such, we have that
$\mathrm{VS}_{1}(Z)\leq\frac{T}{n}D_{1}(Z,\frac{1}{n})$ (57)
proving the theorem.
Appendix C Notes on previous versions and alternative proofs
C.1 Proof that $VS\geq LCR$ for $q\in\{2,3\}$
In a previous version of this paper, 1 was left as a conjecture except for $q\in{2,3,+\infty}$, and $q\in[-\infty,0]$ when $Z$ is full rank. The proofs for $q=+\infty$ and for $q\in[-\infty,0]$ were as currently presented in Section III. The previous proofs for $q\in\{2,3\}$ are presented below
Proof of 1 in the special cases of $q=2$ and $q=3$.
At $q=2$,
[…]
$\frac{n^{2}}{\sum_{i,j}Z^{2}_{i,j}}\geq\frac{n^{2}}{\sum_{i,j}Z_{i,j}}$ (59)
For $q=3$, we have that
[…]
This gives
[…]
proving the $q=3$ case. ∎
C.2 Hoffman 1966 and a proof for integer $q\geq 2$
A previous version of this paper conjectured the following result, noting that it would help in proving 3 (which was also stated as a conjecture).
Theorem 2 (Hoffman 1966).
Denote by $\mathbf{1}$ the $n$ dimensional vector each of whose entries are $1$, Let $M$ be any $n\times n$ symmetric matrix with non-negative entries, and denote by $r=M\mathbf{1}$ (i.e. $r_{i}=\sum_{j}M_{ij}$) the row-sums of $M$. Then for any integer $\alpha\geq 0$
[…]
Since then we have learned that this result was proved by Hoffman in [19] following up on earlier work by London in [30], which had proved the special cases of $n\leq 3$ for all integers $\alpha\geq 1$, and $5\geq\alpha\geq 1$ for arbitrary $n>0$. This result leads to an alternative proof of 3 for integer $q\geq 1$ (and thus a proof of 1 for integer $q\geq 2$), which is essentially what we had in mind in the earlier version of this work. This proof is as follows:
Proof of 3 in the special case of integer $q\geq 1$.
Consider that
[…]
since $Z$ is symmetric, $0\leq Z_{i,j}\leq 1$ for each $i,j$, and $Z^{q-1}$ is non-negative (since it is the product of q1q-1 non-negative matrices). But since $Z$ is symmetric and non-negative, and by assumption $q-1\geq 0$ is an integer, by 2
[…]
C.3 Python code to search for counterexamples
In a previous version of this paper, we discussed a numerical search for counterexamples to 1 (partly inspred by [13]). The code used for that search is provided below.
Note this does not search over the whole space of non-negative PSD similarity matrices. Rather, by construction, it only searches over completely positive matrices. For $n$ by $n$ matrices with $n\leq 4$, all non-negative PSD matrices are completely positive, but for $n\geq 5$ this is not the case [32]. Given the improvements in what we have proven over that version, we no longer see a need for such a numerical counterexample search; the code is provided only for archival reasons.
@@ C.3 Python code to search for counterexamples
# keep only the support
evals = np.array(
[e for e in evals if e>0]
@@
# Compute vendi score
ret = vendi_score(Z, q)
# Compute and subtract
# uniform diversity
ret -= uniform_div(Z, q)
return ret
@@
# map from reals to positive
# reals, reshape
@@
combos = [{
’q’: q,
’n’: n,
’r’: r,
’rep’: rep,
’scale’: s,}
for q in qlist for n in sizes for s in scales
for rep in replicates
@@
rows = pool.map(
map_func,
combos[ix_start:ix_end],
@@
results_df.to_pickle(outfile)